# Positive and Unlabeled Data: Model, Estimation, Inference, and Classification

### v2

@@ 3.1 Empirical Likelihood Method
where $N\equiv n+m$. In addition, $(\alpha_{1},\alpha_{2},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2})$ and $p_{s}({\bf x}|y=1)$ must satisfy the conditions in (5), which can rewritten as
@@
We propose to use the EL method (Owen, 1990; Qin and Lawless, 1994; Owen, 2001) to estimate parameters. The EL method allows us to estimate parameters without assuming a specific distribution assumption on $p_{s}({\bf x}|y=1)$. Specifically, we denote $p_{i}=p_{s}({\bf x}_{i}|y=1)$ for $i=1,\dots,N$ and treat them as parameters. Then the log-EL function is given by
$\begin{split}&\ell_{\text{EL}}(\alpha_{1},\alpha_{2},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2},\pi,p_{1},\dots,p_{N})\\ =&\sum_{i=1}^{N}{\rm log}(p_{i})+\sum_{j=1}^{m}{\rm log}\left\{\pi\exp\left(\alpha_{1}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}_{1}\right)+(1-\pi)\exp\left(\alpha_{2}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}\right)\right\}.\end{split}$ (9)
Noe that the feasible values of $p_{i},i=1,\dots,N$ must satisfy following conditions
@@
The explicit form of $\widehat{\boldsymbol{\theta}}$ is generally unknown. In the next subsection, we present an EM-algorithm to numerically compute $\widehat{\boldsymbol{\theta}}$.
@@ 3.2 Expectation-Maximization algorithm
$\begin{split}\ell_{\mathrm{EL}}^{C}({\boldsymbol{\Theta}})=&\sum_{i=1}^{N}{\rm log}\left(p_{i}\right)+\sum_{j=1}^{m}\left\{y_{n+j}(\alpha_{1}+{\bf x}_{n+j}^{\rm T}{\boldsymbol{\beta}}_{1})+(1-y_{n+j})(\alpha_{2}+{\bf x}_{n+j}^{\rm T}{\boldsymbol{\beta}}_{2})\right\}\\ &+\sum_{j=1}^{m}\left\{y_{n+j}{\rm log}(\pi)+(1-y_{n+j}){\rm log}(1-\pi)\right\},\end{split}$
where ${\boldsymbol{\Theta}}=(\alpha_{1},\alpha_{2},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2},\pi,p_{1},\dots,p_{N})$. The EM algorithm is based on the complete log-EL.
@@
    $\begin{split}Q^{(r)}(\alpha_{1}^{\ast},\alpha_{2}^{\ast},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2})=&-\sum_{i=1}^{N}{\rm log}\left\{{1+\sum_{t=1}^{2}\exp(\alpha_{t}^{\ast}+{\bf x}_{i}^{\rm T}{\boldsymbol{\beta}}_{t})}\right\}\\ &+\sum_{j=1}^{m}\left\{\omega_{n+j}^{(r)}\left(\alpha_{1}^{\ast}+{\bf x}_{n+j}^{\rm T}{\boldsymbol{\beta}}_{1}\right)+(1-\omega_{n+j}^{(r)})(\alpha_{2}^{\ast}+{\bf x}_{n+j}^{\rm T}{\boldsymbol{\beta}}_{2})\right\},\\ \end{split}$ (13)
@@
    This objective function is proportional to a weighted log-likelihood function of a multinomial logistic regression model with three classes. Thus, in practice, we can easily calculate $(\alpha_{1}^{\ast(r)},\alpha_{2}^{\ast(r)},{\boldsymbol{\beta}}_{1}^{(r)},{\boldsymbol{\beta}}_{2}^{(r)})$ by fitting a multinomial logistic regression, which can be done by most software.
@@
Summarizing the above steps leads to Algorithm 1. We call this parameter estimation method the double exponential tilting-empirical likelihood (DET-EL) method.
Data: Source data $\left\{{\bf x}_{i},y_{i}=1\right\}_{i=1}^{n}$; target data $\left\{{\bf x}_{i}\right\}_{i=n+1}^{N}$.
Result: Estimates of ${\boldsymbol{\Theta}}$.
initialization;
while not converged do
       E-step: For $j=1,\dots,m$, compute:
$\omega_{n+j}^{(r)}=\frac{\pi^{(r-1)}\exp(\alpha_{1}^{(r-1)}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}_{1}^{(r-1)})}{\pi^{(r-1)}\exp(\alpha_{1}^{(r-1)}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}_{1}^{(r-1)})+(1-\pi^{(r-1)})\exp(\alpha_{2}^{(r-1)}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}_{2}^{(r-1)})}.$
M-step:
  1. 1.
    Update $\pi$ by $\pi^{(r)}=m^{-1}\sum_{j=1}^{m}\omega_{n+j}^{(r)};$
  2. 2.
    Update $(\alpha_{1}^{\ast},\alpha_{2}^{\ast},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2})$ by
        $(\alpha_{1}^{\ast(r)},\alpha_{2}^{\ast(r)},{\boldsymbol{\beta}}_{1}^{(r)},{\boldsymbol{\beta}}_{2}^{(r)})=\operatorname*{arg\,max}_{\alpha_{1}^{\ast},\alpha_{2}^{\ast},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2}}Q^{(r)}(\alpha_{1}^{\ast},\alpha_{2}^{\ast},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2}),$
    where $Q^{(r)}(\alpha_{1}^{\ast},\alpha_{2}^{\ast},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2})$ is given in (13);
  • 3.
    Update $\alpha_{1}$ and $\alpha_{2}$ by
    $\alpha_{1}^{(r)}=\alpha_{1}^{\ast(r)}-{\rm log}\left(\frac{\sum_{j=1}^{m}\omega_{n+j}^{(r)}}{n}\right),\quad\alpha_{2}^{(r)}=\alpha_{2}^{\ast(r)}-{\rm log}\left(\frac{\sum_{j=1}^{m}(1-\omega_{n+j}^{(r)})}{n}\right);$
  • 4.
    For $i=1,\ldots,N$, update $p_{i}$’s by
    $p_{i}^{(r)}=n^{-1}\left\{1+\exp(\alpha_{1}^{\ast(r)}+{\bf x}_{i}^{\rm T}{\boldsymbol{\beta}}_{1}^{(r)})+\exp(\alpha_{2}^{\ast(r)}+{\bf x}_{i}^{\rm T}{\boldsymbol{\beta}}_{2}^{(r)})\right\}^{-1}.$
  • end while
    Determine label switching and output the estimates.
    Algorithm 1 DET-EL Method for Parameter Estimation
@@
    $\ell_{\text{EL}}({\boldsymbol{\Theta}})=\sum_{i=1}^{n}{\rm log}(p_{i})+\sum_{j=1}^{m}{\rm log}\left\{\pi p_{n+j}\exp\left(\alpha_{1}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}_{1}\right)+(1-\pi)p_{n+j}\exp\left(\alpha_{2}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}\right)\right\},$
    which, together with (10), implies that $\ell_{\text{EL}}({\boldsymbol{\Theta}})\leq 0$. With this result, Proposition 1 guarantees that the EM algorithm eventually converges to at least a local maximum for the given initial value ${\boldsymbol{\Theta}}^{(0)}$. We recommend using multiple initial values to explore the likelihood function to ensure achieving the global maximum. Second, in practice, we may stop the algorithm when the increment in the log-EL after an iteration is no greater than, say, 1e-6.
@@ 4.1 Statistical Inference
    In this section, we first provide the asymptotic distribution of the MELE $\widehat{\boldsymbol{\theta}}$. Then, we focus on two specific statistical inference tasks: the first one is the goodness-of-fit test of the SETM as opposed to the DETM; the second one is inference on the mixture proportion $\pi$.
    We need some notation for developing the asymptotic distribution for $\widehat{\boldsymbol{\theta}}$. Let $(\widehat{\lambda}_{1},\widehat{\lambda}_{2})$ be the Lagrange multipliers corresponding to $\widehat{\boldsymbol{\theta}}$, i.e., $(\widehat{\lambda}_{1},\widehat{\lambda}_{2})$ is the solution of (11) with ${\boldsymbol{\theta}}$ being replaced by $\widehat{\boldsymbol{\theta}}$. Throughout the paper, we assume that the total sample size $N=n+m\to\infty$ and $m/N\to c$ for some constant $c\in(0,1)$. This assumption indicates that $n$ and $m$ go to infinity at the same rate. For simplicity and presentation convenience, we write $c=m/N$ and assume it is a constant, which does not affect our technical development. The following theorem establishes the joint asymptotic normality of $(\widehat{\lambda}_{1},\widehat{\lambda}_{2})$ and $\widehat{\boldsymbol{\theta}}$, which implies the asymptotic normality of $\widehat{\boldsymbol{\theta}}$. The proofs for results in this section will be given in the supplementary materials.
@@
    Suppose the DETM and regularity conditions in Appendix are satisfied. Let ${\boldsymbol{\theta}}^{o}=(\alpha_{1}^{o},\alpha_{2}^{o},{\boldsymbol{\beta}}_{1}^{o},{\boldsymbol{\beta}}_{2}^{o},\pi^{o})$ be the true value of ${\boldsymbol{\theta}}$. As $N$ goes to infinity,
@@
    in distribution, where ${\boldsymbol{\Sigma}}$ is given in (16) in Appendix.
    For SCAR PU data, using the DETM might not be the best approach as SETM could improve computational and estimation efficiency in such cases. To help us decide whether the SETM can provide a good fit to the PU data, we can use the profile log-EL function in (12) to construct the goodness-of-fit test of the SETM as opposed to the DETM. Specifically, the empirical likelihood ratio (ELR) test for testing SETM vs. DETM can be formulated as the following null hypothesis $H_{0}:\alpha_{1}=0$ and ${\boldsymbol{\beta}}_{1}={\bf 0}$ with test statistic given by
@@
    where $\widetilde{\boldsymbol{\theta}}$ is the MELE of ${\boldsymbol{\theta}}$ under the SETM with $\alpha_{1}=0$ and ${\boldsymbol{\beta}}_{1}={\bf 0}$, and $\widehat{\boldsymbol{\theta}}$ is the MELE under the DETM. Note that $\widetilde{\boldsymbol{\theta}}$ can be obtained by slightly modifying Algorithm 1: we only need to fix $\alpha_{1}=0$ and ${\boldsymbol{\beta}}_{1}={\bf 0}$ in the E-step and M-step of Algorithm 1.
@@
    where $\widehat{\boldsymbol{\theta}}_{\pi}$ is the MELE of ${\boldsymbol{\theta}}$ with $\pi$ being fixed. We can derive $\widehat{\boldsymbol{\theta}}_{\pi}$ as if we know the value of $\pi$ in Algorithm 1: we only need to fix $\pi$ in the E-step and M-step of Algorithm 1.
@@
    Assume the same conditions as Theorem 2. We have
@@ 4.2 Approximation of the Bayes Classifier
    Assume the same conditions as Theorem 2. We have
@@ 5 Numerical Studies
    In this section, we first present an ablation study to evaluate the proposed method. Then, we apply our method to a real data application to predict the selling price of mobile phones. We conducted the numerical study on a MacBook with an M3 Pro chip and 36 GB of RAM, using R version 4.4.0 (R Core Team,, 2024), and our custom package PUEM given in the supplementary materials. Some benchmark methods were implemented in Python version 3.11.7 because the authors only provide Python source code.
    5.1 Ablation Study
    We conduct an ablation study to assess the performance of our proposed framework across various tasks. We are interested in the following tasks: 1) Conducting a goodness-of-fit test between SETM and DETM, focusing on the type-I error and power of the test procedure. 2) Evaluating the mean square error (MSE) of the MELE of $\pi$. 3) Determining the coverage accuracy of the confidence intervals of $\pi$. 4) Assessing the classification accuracy using the approximated Bayes classifier.
    Detection of SCAR Data
    Letting ${\bf 0}_{p}$ and ${\bf 1}_{p}$ column vectors of 00’s and $1$’s with dimension $p$, respectively, we generated data with the following multivariate normal distributions: 1) Source positive distribution: $p_{s}({\bf x}|y=1)\sim\mathrm{MVN}(\boldsymbol{\mu}_{1},\mathbf{I})$, where $\boldsymbol{\mu}_{1}={\bf 0}_{15}\in{\bf R}^{15}$. 2) Target negative distribution: $p_{t}({\bf x}|y=0)\sim\mathrm{MVN}(\boldsymbol{\mu}_{2},\mathbf{I})$, where $\boldsymbol{\mu}_{2}={\bf 1}_{15}\in{\bf R}^{15}$. 3) Target positive distribution: $p_{t}({\bf x}|y=1)\sim\mathrm{MVN}(\boldsymbol{\mu}_{3},\mathbf{I})$, where we can choose different levels for $\boldsymbol{\mu}_{3}$: $\boldsymbol{\mu}_{3}\in\left\{{\bf 0}_{15},(1~{}{\bf 0}_{14}^{\rm T})^{\rm T},\dots,({\bf 1}_{7}^{\rm T}~{}{\bf 0}_{8}^{\rm T})^{\rm T}\right\}.$ We shall note that when $\boldsymbol{\mu}_{3}={\bf 0}_{15}$, the PU data are SCAR; otherwise, the data are SAR. We performed the goodness-of-fit test for SETM vs. DETM, where the null hypothesis is that the PU data are SCAR and follow the SETM, which can be translated to $H_{0}:\alpha_{1}=0\mbox{ and }{\boldsymbol{\beta}}_{1}={\bf 0}_{15}$.
    In the experiment, we set the sample size of the source data equal to that of the target data ($n=m$) and fixed the proportion of positive data on the target at 75% ($\pi=0.75$). The levels of $n$ were $1000,2000,3000,4000,$ and $5000$. Each factor level combination was repeated 500 times. Table 1 summarizes the rejection rates of the goodness-of-fit test at the 5% significance level.

### v3

@@ 3.1 Empirical Likelihood Method
where $N\equiv n+m$. In addition, $(\alpha_{1},\alpha_{2},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2})$ and $p_{s}({\bf x}|y=1)$ must satisfy the conditions in (5), which can be rewritten as
@@
We propose to use the EL method (Qin and Lawless, 1994; Owen, 2001) to estimate parameters. The EL method allows us to estimate parameters without assuming a specific distribution assumption on $p_{s}({\bf x}|y=1)$. Specifically, we denote $p_{i}=p_{s}({\bf x}_{i}|y=1)$ for $i=1,\dots,N$ and treat them as parameters. Then the log-EL function is given by
$\begin{split}&\ell_{\text{EL}}(\alpha_{1},\alpha_{2},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2},\pi,p_{1},\dots,p_{N})\\ =&\sum_{i=1}^{N}{\rm log}(p_{i})+\sum_{j=1}^{m}{\rm log}\left\{\pi\exp\left(\alpha_{1}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}_{1}\right)+(1-\pi)\exp\left(\alpha_{2}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}_{2}\right)\right\}.\end{split}$ (9)
Note that the feasible values of $p_{i},i=1,\dots,N$ must satisfy the following conditions
@@
The explicit form of $\widehat{\boldsymbol{\theta}}$ is generally unknown. In theory, a Newton-type algorithm could be employed to numerically calculate $\widehat{\boldsymbol{\theta}}$. However, because $\ell_{N}({\boldsymbol{\theta}})$ is not concave, the convergence of such an algorithm is not guaranteed. Instead, the following subsection introduces an EM algorithm for numerically computing $\widehat{\boldsymbol{\theta}}$, which offers a theoretical guarantee of convergence.
@@ 3.2 Expectation-Maximization Algorithm
$\begin{split}\ell_{\mathrm{EL}}^{C}({\boldsymbol{\Theta}})=&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{i=1}^{n}{\rm log}\left\{p_{s}({\bf x}_{i}|y=1)\right\}}\\ &{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\sum_{j=1}^{m}{\rm log}\left[\{\pi p_{t}({\bf x}_{n+j}|y=1)\}^{y_{n+j}}\{(1-\pi)p_{t}({\bf x}_{n+j}|y=0)\}^{1-y_{n+j}}\right]}\\ =&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{i=1}^{n}{\rm log}\left\{p_{s}({\bf x}_{i}|y=1)\right\}}\\ &{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\sum_{j=1}^{m}\left[y_{n+j}{\rm log}\left\{\pi p_{t}({\bf x}_{n+j}|y=1)\right\}+(1-y_{n+j}){\rm log}\left\{(1-\pi)p_{t}({\bf x}_{n+j}|y=0)\right\}\right]}\\ =&\sum_{i=1}^{N}{\rm log}\left(p_{i}\right)+\sum_{j=1}^{m}\left\{y_{n+j}(\alpha_{1}+{\bf x}_{n+j}^{\rm T}{\boldsymbol{\beta}}_{1})+(1-y_{n+j})(\alpha_{2}+{\bf x}_{n+j}^{\rm T}{\boldsymbol{\beta}}_{2})\right\}\\ &+\sum_{j=1}^{m}\left\{y_{n+j}{\rm log}(\pi)+(1-y_{n+j}){\rm log}(1-\pi)\right\},\end{split}$
where ${\boldsymbol{\Theta}}=(\alpha_{1},\alpha_{2},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2},\pi,p_{1},\dots,p_{N})$ and we have used the DETM (4) in the last step. The EM algorithm is based on the complete log-EL.
@@
    $\begin{split}Q^{(r)}(\alpha_{1}^{\ast},\alpha_{2}^{\ast},{\boldsymbol{\beta}}_{1},{\boldsymbol{\beta}}_{2})=&-\sum_{i=1}^{N}{\rm log}\left\{{1+\sum_{t=1}^{2}\exp(\alpha_{t}^{\ast}+{\bf x}_{i}^{\rm T}{\boldsymbol{\beta}}_{t})}\right\}\\ &+\sum_{j=1}^{m}\left\{\omega_{n+j}^{(r)}\left(\alpha_{1}^{\ast}+{\bf x}_{n+j}^{\rm T}{\boldsymbol{\beta}}_{1}\right)+(1-\omega_{n+j}^{(r)})(\alpha_{2}^{\ast}+{\bf x}_{n+j}^{\rm T}{\boldsymbol{\beta}}_{2})\right\},\\ \end{split}$
@@
    This objective function in this optimization problem is proportional to a weighted log-likelihood function of a multinomial logistic regression model with three classes. Thus, in practice, we can easily calculate $(\alpha_{1}^{\ast(r)},\alpha_{2}^{\ast(r)},{\boldsymbol{\beta}}_{1}^{(r)},{\boldsymbol{\beta}}_{2}^{(r)})$ by fitting a multinomial logistic regression, which can be done by most software.
@@
A summarization of the previous steps is given in the Section S2 of the supplementary materials.
@@
$\ell_{\text{EL}}({\boldsymbol{\Theta}})=\sum_{i=1}^{n}{\rm log}(p_{i})+\sum_{j=1}^{m}{\rm log}\left\{\pi p_{n+j}\exp\left(\alpha_{1}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}_{1}\right)+(1-\pi)p_{n+j}\exp\left(\alpha_{2}+{\bf x}^{\rm T}_{n+j}{\boldsymbol{\beta}}_{2}\right)\right\},$
which, together with (10), implies that $\ell_{\text{EL}}({\boldsymbol{\Theta}})\leq 0$. With this result, Proposition 1 guarantees that the EM algorithm eventually converges to at least a local maximum for the given initial value ${\boldsymbol{\Theta}}^{(0)}$. Practically, we recommend using multiple initial values to explore the likelihood function to ensure achieving the global maximum. In addition, one can stop the algorithm when the increment in the log-EL after an iteration is no greater than, say, 1e-4.
@@ 4.1 Statistical Inference
In this section, we first provide the asymptotic distribution of the MELE $\widehat{\boldsymbol{\theta}}$. Then, we focus on two specific statistical inference tasks: the first one is the goodness-of-fit test of the SETM as opposed to the DETM; the second one is the inference on the mixture proportion $\pi$.
We need some notation for developing the asymptotic distribution for $\widehat{\boldsymbol{\theta}}$. Let $(\widehat{\lambda}_{1},\widehat{\lambda}_{2})$ be the Lagrange multipliers corresponding to $\widehat{\boldsymbol{\theta}}$, i.e., $(\widehat{\lambda}_{1},\widehat{\lambda}_{2})$ is the solution of (11) with ${\boldsymbol{\theta}}$ being replaced by $\widehat{\boldsymbol{\theta}}$. Throughout the paper, we assume that the total sample size $N=n+m\to\infty$ and $m/N\to c$ for some constant $c\in(0,1)$. This assumption indicates that $n$ and $m$ go to infinity at the same rate. For simplicity and presentation convenience, we write $c=m/N$ and assume it is a constant, which does not affect our technical development. The following theorem establishes the joint asymptotic normality of $(\widehat{\lambda}_{1},\widehat{\lambda}_{2})$ and $\widehat{\boldsymbol{\theta}}$, which implies the asymptotic normality of $\widehat{\boldsymbol{\theta}}$. The proofs for results in this section are given in the supplementary materials.
@@
Suppose the DETM and conditions in Section S4.1 of the supplementary materials are satisfied. Let ${\boldsymbol{\theta}}^{o}=(\alpha_{1}^{o},\alpha_{2}^{o},{\boldsymbol{\beta}}_{1}^{o},{\boldsymbol{\beta}}_{2}^{o},\pi^{o})$ be the true value of ${\boldsymbol{\theta}}$. As $N$ goes to infinity,
@@
in distribution, where ${\boldsymbol{\Sigma}}$ is given in Equation (14) in the supplementary materials.
For CD PU data, using the DETM might not be the best approach as SETM could improve computational and estimation efficiency in such cases. To help us decide whether the SETM can provide a good fit to the PU data, we can use the profile log-EL function in (12) to construct the goodness-of-fit test of the SETM as opposed to the DETM. Specifically, the empirical likelihood ratio (ELR) test for testing SETM vs. DETM can be formulated as the following null hypothesis $H_{0}:\alpha_{1}=0$ and ${\boldsymbol{\beta}}_{1}={\bf 0}$ with test statistic given by
@@
where $\widetilde{\boldsymbol{\theta}}$ is the MELE of ${\boldsymbol{\theta}}$ under the SETM with $\alpha_{1}=0$ and ${\boldsymbol{\beta}}_{1}={\bf 0}$, and $\widehat{\boldsymbol{\theta}}$ is the MELE under the DETM. Note that $\widetilde{\boldsymbol{\theta}}$ can be obtained by slightly modifying the EM algorithm: we only need to fix $\alpha_{1}=0$ and ${\boldsymbol{\beta}}_{1}={\bf 0}$ in the E-step and M-step.
@@
where $\widehat{\boldsymbol{\theta}}_{\pi}$ is the MELE of ${\boldsymbol{\theta}}$ with $\pi$ being fixed. We can derive $\widehat{\boldsymbol{\theta}}_{\pi}$ as if we know the value of $\pi$ in the EM algorithm: we only need to fix $\pi$ in the E-step and M-step.
@@
Assume the same conditions as in Theorem 2. We have
@@ 4.2 Approximation of the Bayes Classifier
Assume the same conditions as in Theorem 2. We have
@@ 5 Numerical Studies
In this section, we first present a simulation study to evaluate the proposed methods. Then, we apply our method to a real data application to predict the selling price of mobile phones. We use our custom R package PUEM in this section, the package is available in the supplementary materials. Some benchmark methods were implemented in Python because the authors only provide Python source code.
5.1 Simulation Study
We conduct a simulation study to assess the performance of our proposed framework across various tasks. We are interested in the following tasks: 1) Conducting a goodness-of-fit test between SETM and DETM, focusing on the type-I error and power of the test procedure. 2) Evaluating the root mean square error (RMSE) of the MELE of $\pi$. 3) Determining the coverage accuracy of the confidence intervals of $\pi$. 4) Assessing the classification accuracy using the approximated Bayes classifier.
Detection of CD Data
Let ${\bf 0}_{p}$ and ${\bf 1}_{p}$ denote the p-dimensional column vectors of zeros and ones, respectively. We generate data from the following multivariate normal distributions: 1) Source positive distribution: $p_{s}({\bf x}|y=1)\sim\mathrm{MVN}(\boldsymbol{\mu}_{1},\mathbf{I})$, where $\boldsymbol{\mu}_{1}={\bf 0}_{15}\in\mathbb{R}^{15}$ and ${\bf I}$ is the identity matrix. 2) Target negative distribution: $p_{t}({\bf x}|y=0)\sim\mathrm{MVN}(\boldsymbol{\mu}_{2},\mathbf{I})$, where $\boldsymbol{\mu}_{2}={\bf 1}_{15}\in\mathbb{R}^{15}$. 3) Target positive distribution: $p_{t}({\bf x}|y=1)\sim\mathrm{MVN}(\boldsymbol{\mu}_{3},\mathbf{I})$, where we can choose different levels for $\boldsymbol{\mu}_{3}\in\left\{{\bf 0}_{15},(1~{}{\bf 0}_{14}^{\rm T})^{\rm T},\dots,({\bf 1}_{7}^{\rm T}~{}{\bf 0}_{8}^{\rm T})^{\rm T}\right\}.$ It should be noted that when $\boldsymbol{\mu}_{3}={\bf 0}_{15}$, the PU data are CD; otherwise, the data are DD. We perform the goodness-of-fit test for SETM vs. DETM, where the null hypothesis is that the PU data are CD and follow the SETM, which can be translated to $H_{0}:\alpha_{1}=0\mbox{ and }{\boldsymbol{\beta}}_{1}={\bf 0}_{15}$.
In the experiment, we set the sample size of the source data equal to that of the target data ($n=m$) and fix the proportion of positive data on the target at 75% ($\pi=0.75$). The levels of $n$ are $1000,2000,3000,4000,$ and $5000$. Each factor level combination is repeated 1000 times. Table 1 summarizes the rejection rates of the goodness-of-fit test at the 5% significance level.